# cuSZ-$i$: High-Ratio Scientific Lossy Compression on GPUs with Optimized Multi-Level Interpolation

### v4

@@ VI-B Synergy of Lossless Modules: Huffman + Bitcomp-Lossless
Though the Huffman-coding-only practice leads to a balanced throughput and compression ratio in average cases, its compression is bottlenecked given the fact that each input quant-code must be represented by at least one bit; At the same time, it can still present redundancy of repeated patterns after encoding (e.g., continuous 0x00 bytes). Nevertheless, sophisticated dictionary-based encoders are either limited in throughput (e.g., GPU-LZ [zhang2023gpulz]) or compression ratio on GPU. Though previous work proposed several de-redundancy lossless encoding schemes encoding [cuszplus2021, cuszp, FZGPU] (alternative to Huffman encoding) that can surpass the one-bit-per-element limitation in extremely compressible cases, they still have unsatisfactory compression ratios in many cases, as evidenced in Fig. 7. Therefore, we propose to synergize Huffman encoding and the subsequent repeated pattern-canceling encoding scheme, which could result in a huge gain in the compression ratio with minimal throughput overhead because this additional scheme only takes inputs of reduced sizes after Huffman encoding. To fit our chart of high ratio and high throughput, we selected Bitcomp-Lossless [bitcomp] from NVIDIA, a performance-oriented encoder on GPU, after trial and error over a large variety of existing GPU-based bitstream lossless encoders, We are aware of the proprietary nature of NVIDIA’s Bitcomp; however, it only serves for the demonstration purposes for the effectively utilizing the created high compressibility from G-Interp on GPU. In §VII, we will present how Bitcomp-Lossless helps to greatly remove the redundancy of Huffman-encoded quant-codes generated by cuSZ-$i$ data predictor with negligible computational overhead.
@@ VII-A Experimental Setup
Platforms
We evaluate on §VI-A tweak§VI-B encoding synergy1 NVIDIA A100 GPUs from ALCF-ThetaGPU [testbed-thetagpu] and Purdue-Anvil [testbed-anvil] for profiling and Globus test and §VI-A tweak§VI-B encoding synergy2 NVIDIA A40 GPU on ANL-JLSE [testbed-jlse] for profiling. More details are in TABLE I.
Baselines
@@
compute 19.49 TFLOPS 37.42 TFLOPS
CUDA/driver 11.4/470.161.03 11.6/530.30.02 11.8/545.23.06
@@
10 files dim: $zyx=512\times 512\times 512$ total: 5 GB
@@
7 files dim: $zyx=256\times 384\times 384$ total: 1 GB
@@
6 files dim: $zyx=512\times 512\times 512$ total: 3.1 GB
@@
1 files dim: $zyx=288\times 115\times 69\times 69$ total: 612 MB
@@
37 files dim: $zyx=449\times 449\times 235$ total: 6.5 GB
@@
11 files dim: $zyx=500\times 500\times 500$ total: 5.1 GB
@@ VII-B Evaluation Metrics
Fixed-error-bound compression ratio (CR)
@@
Rate-distortion graph
@@
Data transfer time
@@ VII-C Evaluation Results and Analysis
$\blacksquare$ C.1)  Compression Ratios
We compress the datasets under common fixed error bounds and list all the compression ratios in TABLE III (no results for cuZFP as it does not support absolute error-bounding, and no results for cuSZx on dataset Nyx for its runtime errors). TABLE III presents the two-fold achievement of our novel pipeline. §VI-A tweak§VI-B encoding synergy1 In the left half, we present the compression ratios without an extra lossless module; even so, cuSZ-$i$ outperforms other compressors in 14 out of 18 cases, exhibiting 10% to 30% advantages over the second-best ones (column 6). §VI-A tweak§VI-B encoding synergy2 In the right half, we examine the full pipeline integrating Bitcomp-Lossless. For a fair comparison, we apply Bitcomp-Lossless to all other compressors’ outputs. First, the full pipeline sees a significant gain in compression ratio from the one without Bitcomp-Lossless (i.e., comparing column v with 5). Second, cuSZ-$i$ achieves the highest compression ratios in all 18 cases. With the expanded advantage over the second-best with Bitcomp-Lossless enabled (column vi), cuSZ-$i$ tops at 476%. These note the more profound compressibility is created by G-Interp, and is more attuned to the additional pass of lossless encoding than any other compressors. Last, it is worth noticing that, for all results in TABLE III, the decompression PSNR of cuSZ-$i$ is also much higher than other compressors. If we compare the compression ratio under the same PSNR instead of the same error bound, the improvements from cuSZ-$i$ will be amplified (to be detailed in §C.2).
@@
$\blacksquare$ C.2)  Compression Rate-Distortion
Figure 7: (Bit rate)-PSNR graphs of compressors on six datasets. A curve positioned toward upperleft indicates the advantageous performance in both better quality and compression ratios.
To separately exhibit both the contributions of interpolation-based data predictor and the Bitcomp-Lossless to the compression rate-distortion, we present the compression rate-distortion of the compressors in 2 parallel series: one without and one with Bitcomp-Lossless. Fig. 7 displays the compression bit rate-PSNR curves on 6 datasets, without and with Bitcomp-Lossless integrated, respectively. When no extra lossless module is appended, cuSZ-$i$ has already achieved the best rate-distortion, attributed to the effective G-Interp data predictor. On dataset JHTDB, under a PSNR of 70 dB, or on dataset QMCPACK, under a PSNR of 80 dB, cuSZ-$i$ can have around 60% to 80% compression ratio advantages over the second-best cuZFP/cuSZ. In Bitcomp-Lossless-integrated results, cuSZ-$i$ has compression ratios far better than any other, achieving roughly 100% to 500% compression ratio advantages under the same PSNRs in low-bit-rate cases. In high-bit-rate cases, it also promotes considerable reduction rates. The great extent of compression improvement by cuSZ-$i$ originates from its high adaptability to Bitcomp-Lossless since its G-Interp results in highly compressible error quant-codes. As a reference baseline in Fig. 7, we also plot the rate-distortion of QoZ [qoz], which is the state-of-the-art interpolation-based lossy compressor on the CPU platform and has a similar predictor design with cuSZ-$i$. Comparing the rate-distortion curves of cuSZ-$i$ and QoZ, we can make a conclusion that the compression rate-distortion of cuSZ-$i$ (with Bitcomp-Lossless) has been quite close to the best CPU interpolation-based compressors with far better throughputs. Note that CPU-based QoZ has a better compression ratio than cuSZ-$i$ due to larger interpolation blocks and more effective lossless modules.
$\blacksquare$ C.3)  Case Study of Decompression Visualization
(a) JHTDB: the original vs the reconstructed.
(b) S3D: the original vs the reconstructed.
Figure 8: Showcases of visualized decompression data. For each snapshot, we aligned the CR (with Bitcomp-Lossless) from different compressors.
To further verify the high compression quality of cuSZ-$i$, we visualize certain data decompression snapshots from the datasets. Due to space limits, we selected JHTDB and S3D datasets. Fig. 8 shows the original and decompression visualization of two data snapshots by six compressors, together with a note of the compression ratios, error bounds, and PSNRs. For each set of visualizations on the same snapshot, we align the compression ratios to the same value (e.g., $\sim$ 27 for JHTDB).
@@
$\blacksquare$ C.4)  Compression Throughputs
@@
$\blacksquare$ C.5)  Case Study: Distributed Lossy Data Transmission
@@
In Fig. 10, we show the data transfer cost in time and decompression data PSNR for this task. Specifically, the source machine is ALCF Theta-GPU, and the destination machine is Purdue RCAC Anvil. Since the local I/O time is irrelevant to this scenario, we compute the overall time using only the data compression/decompression time and the distributed compressed data transfer time. The data transfer between 2 machines is managed by Globus [globus2] with a bandwidth of approximately 1 GB/s. Moreover, the full pipeline with Bitcomp-Lossless is applied to all cases to further reduce data transfer costs. As Fig. 10 indicates, cuSZ-$i$ has the best-in-class time cost on all high-quality data transfer cases (PSNR $\leq$ 70 dB). cuSZ-$i$ is also very competitive on specific data transfer tasks with low-quality data. For dataset QMCPACK, cuSZ-$i$ can reduce around 30% of the time cost of the second-best cuSZ for transferring at a PSNR of 90 dB. For dataset S3D, it can reduce around 40% of the time cost of the second-best cuZFP when transferring at a PSNR of 80 dB. In short, it is evident that the high compression ratio and quality of cuSZ-$i$ have gracefully compensated for its speed degradation, exhibiting great potential for a great variety of real-world use cases.
@@ Acknowledgments
This research was supported by the Exascale Computing Project (ECP), Project Number: 17-SC-20-SC, a collaborative effort of two DOE organizations—the Office of Science and the National Nuclear Security Administration, responsible for the planning and preparation of a capable exascale ecosystem, including software, applications, hardware, advanced system engineering and early testbed platforms, to support the nation’s exascale computing imperative. The material was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under contracts DE-AC02-06CH11357. This work was also supported by the National Science Foundation under Grants 2003709, 2303064, 2104023, 2247080, 2247060, 2312673, 2311875, and 2311876.

### v5

@@ VI-B Synergy of Lossless Modules: Huffman + Bitcomp-lossless
Though the Huffman-coding-only practice balances throughput and compression ratio in average cases, its compression is obstructed given that each input quant-code must be represented by at least one bit; At the same time, it can still present redundancy of repeated patterns after encoding (e.g., continuous 0x00 bytes). Nevertheless, sophisticated dictionary-based encoders are either limited in throughput (e.g., GPU-LZ [zhang2023gpulz]) or compression ratio on GPU. Though previous work proposed several de-redundancy lossless encoding schemes encoding [cuszplus2021, cuszp, FZGPU] (alternative to Huffman encoding) that can surpass the one-bit-per-element limitation in extremely compressible cases, they still have unsatisfactory compression ratios in many cases, as evidenced in Fig.7(a). Therefore, we propose synergizing Huffman encoding and the subsequent repeated pattern-canceling encoding scheme, which could result in a huge gain in the compression ratio with minimal throughput overhead because this additional scheme only takes inputs of reduced sizes after Huffman encoding. To fit our chart of high ratio and high throughput, we selected Bitcomp-lossless [bitcomp] from NVIDIA, a performance-oriented encoder on GPU, after trial and error over a large variety of existing GPU-based bitstream lossless encoders, We are aware of the proprietary nature of NVIDIA’s Bitcomp; however, it only serves for demonstration purposes for effectively utilizing the created high compressibility from G-Interp on GPU. In §VII, we will present how Bitcomp-lossless helps to greatly remove the redundancy of Huffman-encoded quant-codes generated by cuSZ-$i$ data predictor with negligible computational overhead.
@@ VII-A Experimental Setup
$\square$ Platforms.
We evaluate on VI-B1 NVIDIA A100 GPUs from ALCF-ThetaGPU [testbed-thetagpu] and Purdue-Anvil [testbed-anvil] for profiling and Globus test and VI-B2 NVIDIA A40 GPU on ANL-JLSE [testbed-jlse] for profiling. More details are in TABLE I.
$\square$ Baselines.
@@
compute (FP32)
19.49 TFLOPS 37.42 TFLOPS
CUDA version 11.4 11.6 11.8
driver version 470.161.03 530.30.02 545.23.06
@@
10 files dim: $512_{z}\times 512_{y}\times 512_{x}$ total: 5 GB
@@
7 files dim: $256_{z}\!\times 384_{y}\!\times 384_{x}$ total: 1 GB
@@
6 files dim: $512_{z}\!\times 512_{y}\!\times 512_{x}$ total: 3.1 GB
@@
1 files dim: $(288\times 115)_{z}\!\times 69_{y}\!\times 69_{x}$ total: 612 MB
@@
37 files dim: $449_{z}\!\times 449_{y}\!\times 235_{x}$ total: 6.5 GB
@@
11 files dim: $500_{z}\!\times 500_{y}\!\times 500_{x}$ total: 5.1 GB
@@ VII-B Evaluation Metrics
$\square$ Fixed-error-bound compression ratio (CR).
@@
$\square$ Rate-distortion graph.
@@
$\square$ Data transfer time.
@@ VII-C Evaluation Results and Analysis
$\blacksquare$ C.1)  Compression Ratios.
We compress the datasets under fixed error bounds that are commonly used and list all the results in TABLE III (no results for cuZFP as it does not support absolute error-bounding, and no results for cuSZx on dataset Nyx for its runtime errors). TABLE III presents the two-fold achievement of our novel pipeline. VI-B1 In the left half, we present the compression ratios without an extra lossless module; even so, cuSZ-$i$ outperforms others in 14 of 18 cases, exhibiting 10% to 30% advantages over the second-bests (column 6). VI-B2 In the right half, we examine the full pipeline integrating Bitcomp-lossless. For fairness, we apply Bitcomp-lossless to all other compressors’ outputs. First, the full pipeline sees a significant gain in compression ratio from the one without Bitcomp-lossless (i.e., comparing column v with 5). Second, cuSZ-$i$ achieves the highest compression ratios in all 18 cases. cuSZ-$i$ expands advantage over the second-best with Bitcomp-lossless enabled (column vi) and tops at 476%. These note that G-Interp creates more profound compressibility and is more attuned to the additional pass of lossless encoding than any other compressors. Last, it is worth noticing that, for all results in TABLE III, the decompression PSNR of cuSZ-$i$ is also much higher than other compressors. If we compare the compression ratio under the same PSNR instead of the same error bound, the improvements from cuSZ-$i$ will be amplified (to be detailed in §C.2).
@@
$\blacksquare$ C.2)  Compression Rate-Distortion.
(a) (Bit rate)-PSNR graphs from compression results without (top row) and with Bitcomp-lossless (bottom row) encoding pass.
VI-BRefer to captionfixed PSNR,smaller bit rate
(b) (Bit rate)-PSNR leftward change due to the fixed PSNR and the smaller bit rate from the extra pass of lossless encoding (Bitcomp-lossless).
Figure 7: Compression (bit rate)-PSNR graphs on six datasets. A curve toward upperleft indicates both advantageous quality and compression ratios.
Both Interpolation-based data predictor and Bitcomp-lossless contribute to better compression rate-distortion trait. To exhibit the separate contributions, we present the rate-PSNR curves on 6 datasets in 2 parallel series in Fig.7(a): without and with Bitcomp-lossless. When no extra lossless module is appended, cuSZ-$i$ has already achieved the best rate-distortion, attributed to the effective G-Interp data predictor. On dataset JHTDB under a PSNR of 70 dB, or on dataset QMCPACK under a PSNR of 80 dB, cuSZ-$i$ shows 60% to 80% advantages in compression ratio over the second-best cuZFP/cuSZ. In the with-Bitcomp results, cuSZ-$i$ has far better compression ratios than any other, achieving roughly 100% to 500% advantages under the same PSNRs in low-bit-rate cases. In high-bit-rate cases, it also promotes considerable reduction rates. The great extent of compression boost by cuSZ-$i$ originates from its high adaptability to Bitcomp-lossless since its G-Interp results in highly compressible error quant-codes, which is visualized as the fixed-PSNR bit rate change in auxiliary Fig.7(b). As a reference baseline in Fig.7(a), we also include the rate-distortion of QoZ [qoz], which reflects the latest interpolation-based art on the CPU platform and one design basis of cuSZ-$i$. We conclude that 1) CPU-based QoZ still features a better compression ratio than cuSZ-$i$ due to larger interpolation blocks and more effective lossless modules, but cuSZ-$i$ features far better throughputs; 2) nevertheless, cuSZ-$i$ (with Bitcomp-lossless) has, for the first time, approximated the best-in-class CPU interpolation-based compressor (i.e., QoZ) in rate-distortion trait.
$\blacksquare$ C.3)  Case Study of Decompression Visualization.
Figure 8: Showcases of visualized decompressed data, with key subregions zoomed in. For each snapshot, we align the (with-Bitcomp) compression ratio for every compressor with a fixed values. In addition, cuSZ, FZ-GPU, and cuSZp share the same Lorenzo predictor, exhibiting simiar data visualizations. Hence, they are grouped as the last three.
To further verify the high compression quality of cuSZ-$i$, we visualize certain data decompression snapshots from the datasets. Fig.8 shows the original and decompression visualization of two data snapshots by six compressors, together with a note of the compression ratios, error bounds, and PSNRs. For each set of visualizations on the same snapshot, we align the compression ratios to the same value (e.g., $\sim$ 27 for JHTDB).
@@
$\blacksquare$ C.4)  Compression Throughputs.
@@
$\blacksquare$ C.5)  Case Study: Distributed Lossy Data Transmission.
@@
Fig.10 shows the data-transfer cost in time and decompression data PSNR for this task. Specifically, the source and destination machines are ALCF Theta-GPU and Purdue RCAC Anvil, respectively. Since the local I/O time is irrelevant to this scenario, we compute the overall time only from data compression/decompression time and the data-transfer time of distributed compressed data. Managed by Globus [globus2], the data-transfer bandwidth between the two machines is approximately 1 GB/s. With the full pipeline (including Bitcomp-lossless) fairly applied to all compressors to further reduce data-transfer costs, cuSZ-$i$ has the best-in-class time cost on all high-quality data-transfer cases (PSNR $\geq$ 70 dB), as indicated in Fig.10. cuSZ-$i$ is also very competitive on specific data transfer tasks with low-quality data. For dataset QMCPACK, cuSZ-$i$ can reduce around 30% of the time cost of the second-best cuSZ at a PSNR of 90 dB. For dataset S3D, it can reduce around 40% of the time cost of the second-best cuZFP at a PSNR of 80 dB. Overall, it is evident that the high compression ratio and quality of cuSZ-$i$ have gracefully compensated for its speed degradation, implying significance in a great variety of real-world use cases.
@@ Acknowledgments
This research was supported by the Exascale Computing Project (ECP), Project Number: 17-SC-20-SC, a collaborative effort of two DOE organizations—the Office of Science and the National Nuclear Security Administration, responsible for the planning and preparation of a capable exascale ecosystem, including software, applications, hardware, advanced system engineering and early testbed platforms, to support the nation’s exascale computing imperative. The material was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under contracts DE-AC02-06CH11357. This work was also supported by the National Science Foundation (Grant Nos. 2003709, 2303064, 2104023, 2247080, 2247060, 2312673, 2311875, and 2311876), the National Natural Science Foundation of China (Grant Nos. 62032023 and T2125013), and the Innovation Funding of ICT, CAS (Grant No. E461050).